OpenStudy (anonymous):

What is the lateral area of a cylinder with a radius of 5 ft and a height of 8 ft? Use 3.14 for pie

OpenStudy (shamim):

area of ur cyllinder is \[V=2 \pi rh\]

OpenStudy (shamim):

am i right

OpenStudy (anonymous):

yes

OpenStudy (shamim):

r= 5 ft

OpenStudy (shamim):

h=8 ft

OpenStudy (jack1):

5x 8 = 40 so lateral area =... 80ft^2

OpenStudy (jack1):

?

OpenStudy (anonymous):

so 80 x 40?

OpenStudy (jack1):

no, just 80 sqr ft

OpenStudy (anonymous):

wrong

OpenStudy (shamim):

area=\[2 \times \pi \times 5 \times 8\]

OpenStudy (jack1):

Area = (2x pi x r^2) + (2 x r x ht)

OpenStudy (jack1):

but what is the difference between lateral area and area